a sample of h-like ion is in a particular excited state n2. the electron in it makes back transition up to a lower excited state n1 producing a maximum of 10 different spectral lines. the change in angular momentum of electron corresponding to maximum frequency line is expressed as y(h/2pi)Js then, find the value of y.

Grade:11

1 Answers

Dheeraj
5 years ago
Since total no of spectral line =10
(n1-n2)(n1-n2+1)/2=10
n1-n2=4
Now for max frequency n1-n2=4
     Change in angular momentum=(n1-n2)h/2π
So y=4
Thankyou